OpenStudy (anonymous):

Solve 3x^2-12=-12 by graphing the related function. 1)There is one solution: 24squared 2)There is no solution 3)There are two solutions: +-24squared 4)There are no real number solutions

OpenStudy (anonymous):

@Harkirat can you write this one out too having trouble. thank you!

OpenStudy (anonymous):

you have to solve it by graphing???

OpenStudy (anonymous):

3x^2-12=-12 3x^2-12+12=0 3x^2=0 write it as Y = 3x^2 in your calculator and ask it to plot graph

OpenStudy (anonymous):

yeah i got 0

OpenStudy (anonymous):

yes that is what you will get.....but it does not match any of the answers... check if you typed the equation correctly...

OpenStudy (anonymous):

yeah i typed one wrong i ment "There is one solution: 0"

OpenStudy (anonymous):

ok......then you got yr answer

OpenStudy (anonymous):

yeah i just confused myself!xDthanks